In taking a random sample of 60 automobiles, we find that the odometer readings for the sample have a sample mean of 52,500 and a sample standard deviation of 7,300. At the 95% confidence level, what would be the margin of error for our estimate of the mean odometer reading of the population of automobiles from which this sample was drawn? 6.
Answer #1

Sol:'

margin of error=

T alpha/2,n-1*Sigma/sqrt(n

=t*s/sqrt(n)

Given n=60

n-1=60-1=59

alpha=0.05

alpha/2=0.025

s=sample sd=7300

tcrit59,0.025)==T.INV(0.025;59)=2.001

margin of error=2.001*7300/sqrt(60)=1885.794

ANSWER:1885.794